# Scalable Dynamic Origin-Destination Demand Estimation Enhanced by High-Resolution Satellite Imagery Data

### v1

@@ 4.3 Pittsburgh Downtown network with synthetic data
Following the data generation process described in Section 4.1, we randomly sample 400 links to form the observation link set $A_{obs}$, and introduce stochastic errors to traffic count, speed and density data to simulate sensing errors. The remaining links form the out-of-sample link set $A_{oos}$ for which only density observations with stochastic errors are available. These corrupted observations, including traffic count, speed of observed links and density on all links, are used as input for DODE.
@@
We evaluate the accuracy of estimated traffic conditions separately for observed and unobserved links. This section focuses on analyzing the sensitivity of the DODE framework to varying data conditions, specially the level of sensing error and frequency of imagery-based density snapshots. To reduce the bias introduced by spatial distribution of local sensors, each DODE test is repeated with five different random samples of $A_{obs}$ under each experimental setting.
@@
We first assess the sensitivity of DODE to sensing errors in density observations. As discussed in Section 3.4, density observations may require aggregation when serving as the inputs for DODE and aggregating moving and through vehicles together can introduce sensing errors. Additionally, inaccuracies in class-specific vehicle detection further contribute to noises in sensing results.
@@
To evaluate the impact of such errors, this section considers two levels of sensing errors: (1) ±$\pm 10\%$ with perturbation $\epsilon_{10\%}\sim\text{Unif}(0.9,1.1)$ (2) ±$\pm 20\%$ with perturbation $\epsilon_{20\%}\sim\text{Unif}(0.8,1.2)$. For each error level, 5 DODE runs are conducted using independently sampled observed link sets ($A_{\text{obs}}$) to mitigate the impact related to spatial sampling of local sensors. Consistent with prior sections, $R^{2}$ is used to evaluate model performance. To assess model generalization, we report separate $R^{2}$s for observed and unobserved links to assess the model generalization.
@@
Figure 20 shows the a clear trend that increasing sensor error level from ±$\pm 10\%$ to ±$\pm 20\%$ leads to a noticeable decline in $R^{2}$ values, indicating estimation accuracy is reduced. The flow estimates for unobserved links are relatively more robust, which maintains similar performance under both error level settings, but for density estimation, $R^{2}$ deteriorates with higher error level because sensing errors can reduce the consistency in network-wide density observations. These results demonstrate that the sensing errors in density extraction from imagery data do not have large impact on the local traffic estimation at unobserved locations but will impact overall density estimation itself.
(a) observed car flow estimation
(b) observed truck flow estimation
(c) unobserved car flow estimation
(d) unobserved truck flow estimation
(e) observed car density estimation
(f) observed truck density estimation
@@
Next sensitivity analysis is evaluating the impact of capture frequency of satellite imagery. We fix the same level of sensing errors in density observations as ±$\pm 10\%$ and change density snapshot frequencies in two different levels: (1) density snapshots per 15 min and (2) density snapshots per 30 min. Similarly, 5 DODE runs are conducted in each data setting and separate $R^{2}$s for observed and unobserved links are reported to assess model performance.
(a) observed car flow estimation
(b) observed truck flow estimation
(c) unobserved car flow estimation
(d) unobserved truck flow estimation
(e) observed car density estimation
(f) observed truck density estimation
@@
Figure 21 shows that across all metrics, estimation accuracy, as measured by R-squared, tends to decline as the density snapshot frequency decreases from 15 to 30 minutes This indicates that denser temporal coverage contributes to more accurate calibration.The impact is more pronounced for truck flow estimation than for car flow, as truck volumes are generally lower and thus more sensitive to observational constraints. For density estimations, the 30-min snapshot frequency results in noticeably poorer performance compared to the 15-minute case, likely due to increased uncertainty in the aggregated density observations. Higher snapshot frequency captures more detailed spatiotemporal traffic conditions, which improves estimation accuracy. Similar trends can be observed for car and truck flow estimation on unobserved (out-of-sample) links, suggesting that more frequent density snapshots enhance estimation accuracy for links lacking local sensor data. Nevertheless, estimation accuracy for unobserved links remains lower than that for observed links, reflecting the under-deterministic nature of the DODE problem.
@@ 5 Conclusion
This study proposes a scalable computational graph-based framework to estimate dynamic origin-destination demand by integrating multi-modal data including emerging high-resolution satellite imagery and other conventional mobility data collected from ground-based sensors. Satellite imagery data provides city-wide class-specific traffic density information for both on-road and parking vehicles, which is extracted using computer vision techniques, overcoming the limitations of local ground-based traffic sensor data. A series of numerical experiments using both synthetic and real-world data on networks of various sizes demonstrate that incorporating satellite imagery-derived density can improve estimation performance of the whole network, especially for segments without local sensor coverage. We also conduct sensitivity analysis in terms of sensing error and snapshot frequency in imagery-derived density observations to evaluate the robustness of our model to various data availability conditions.
Future work can be summarized as the following directions: from data perspective, more city-wide satellite imagery data can be collected and used for testing model robustness in real-world applications. For modeling parking behavior, satellite imagery only captures the conditions of curbside spaces and outdoor parking lots. Fusing other parking information such as indoor parking condition and transactions of parking meters will enhance more comprehensive parking choices. From modeling perspective, the computer vision pipeline can be enhanced with advanced lane-level road segmentation features which can differentiate parking and moving vehicles and facilitate more accurate map matching. The computer vision can also be improved for accurate vehicle classification. The proposed model can also be extended to use video-based data which provides traffic information of higher resolution.
@@ Acknowledgment
This research is sponsored by Fujitsu Research of America.
@@
$P^{rs}_{i}$ path set for vehicle class $i$ between $rs$
$f^{rs}_{k,i,t}$ flow on path $k$ for vehicle class $i$ departing at time $t$ between OD pair $rs$
$c^{rs}_{k,i,t}$ travel cost of path $k$ for vehicle class $i$ departing at time $t$ between OD pair $rs$
$p^{rs}_{k,i,t}$ route choice proportion of choosing path $k$ for vehicle class $i$ departing at time $t$ between OD pair $rs$

### v2

@@ 4.3 Pittsburgh Downtown network with synthetic data
Following the data generation process described in Section 4.1, we randomly sample 400 links to form the observation link set $A_{obs}$, and introduce stochastic errors to traffic count and density data to simulate sensing errors. The remaining links form the link set $A_{unobs}$ for which only density observations with stochastic errors are available. These corrupted observations, including traffic counts of observed links and density on all links, are used as input for DODE.
@@
We evaluate the accuracy of estimated traffic conditions separately for observed and unobserved links. This section focuses on analyzing the sensitivity of the DODE framework to varying data conditions, especially the level of sensing error and the frequency of imagery-based density snapshots. To reduce the bias introduced by the spatial distribution of local sensors, each DODE test is repeated with five different random samples of $A_{obs}$ under each experimental setting.
@@
We first assess the sensitivity of DODE to sensing errors in density observations. As discussed in Section 3.4, density observations may require aggregation when serving as inputs for DODE and aggregating moving and through vehicles together can introduce sensing errors. Additionally, inaccuracies in class-specific vehicle detection further contribute to noise in sensing results.
@@
To evaluate the impact of such errors, this section considers two levels of sensing errors: (1) $\pm 10\%$ with perturbation $\epsilon_{10\%}\sim\text{Unif}(0.9,1.1)$ (2) $\pm 20\%$ with perturbation $\epsilon_{20\%}\sim\text{Unif}(0.8,1.2)$. For each error level, 5 DODE runs are conducted using independently sampled observed link sets ($A_{\text{obs}}$) to mitigate the impact related to the spatial sampling of local sensors. Consistent with prior sections, $R^{2}$ is used to evaluate model performance. To assess model generalization, we report separate $R^{2}$ values for observed and unobserved links to evaluate generalization.
@@
Figure 23 shows a clear trend that increasing sensor error level from $\pm 10\%$ to $\pm 20\%$ leads to a noticeable decline in $R^{2}$ values, indicating that estimation accuracy is reduced. The flow estimates for unobserved links are relatively more robust, maintaining similar performance under both error level settings, but for density estimation, $R^{2}$ deteriorates with higher error level because sensing errors can reduce the consistency in network-wide density observations. The perturbation errors also represent the vehicle detection errors and our experiments show that estimation accuracy for cars and trucks (on both observed and unobserved links by local sensors) degrades only modestly under substantial error perturbations. Overall, these results demonstrate that the sensing errors in density extraction from imagery data do not have a large impact on the local traffic estimation at unobserved locations but do impact overall density estimation itself.
Next sensitivity analysis evaluates the impact of the capture frequency of satellite imagery. We fix the level of sensing errors in density observations at $\pm 10\%$ and change density snapshot frequencies to two different levels: (1) one density snapshot per 15 min and (2) one density snapshot per 30 min. Similarly, 5 DODE runs are conducted in each data setting, and separate $R^{2}$ values for observed and unobserved links are reported to assess model performance.
Figure 24 shows that across all metrics, estimation accuracy, as measured by R-squared, tends to decline as the density snapshot frequency decreases from 15 to 30 minutes. This indicates that denser temporal coverage contributes to more accurate calibration.The impact is more pronounced for truck flow estimation than for car flow, as truck volumes are generally lower and thus more sensitive to observational constraints. For density estimation, the 30-min snapshot frequency results in noticeably poorer performance compared to the 15-minute case, likely due to increased uncertainty in the aggregated density observations. Higher snapshot frequency captures more detailed spatiotemporal traffic conditions, which improves estimation accuracy. Similar trends can be observed for car and truck flow estimation on unobserved links, suggesting that more frequent density snapshots enhance estimation accuracy for links lacking local sensor data. Nevertheless, estimation accuracy for unobserved links remains lower than that for observed links, reflecting the underdetermined nature of the DODE problem.
(a) observed car flow
(b) observed truck flow
(c) unobserved car flow
(d) unobserved truck flow
(e) observed car density
(f) observed truck density
@@
(a) observed car flow
(b) observed truck flow
(c) unobserved car flow
(d) unobserved truck flow
(e) observed car density
(f) observed truck density
@@ 5 Conclusion
This study proposes a computational graph-based framework to estimate dynamic origin-destination demand (DODE) by integrating multi-modal data, including emerging high-resolution satellite imagery and conventional traffic data collected from ground-based localized sensors. Satellite imagery data provides city-wide class-specific traffic density information of both on-road and parking vehicles, which can be extracted using computer vision techniques and used as an additional data source in DODE, overcoming the limitations of local traffic sensor data. A series of numerical experiments using both synthetic and real-world data on networks of various sizes demonstrate that incorporating satellite imagery-derived density improves the fit of link density and recovers more plausible dynamic OD patterns, while maintaining satisfactory accuracy for local traffic counts and speeds, especially for segments without local sensor coverage. In the real-world benchmarking, our computational-graph framework also shows faster and more stable convergence and achieves better state estimation accuracy than the PC-SPSA benchmark under the same observation setting, suggesting practical advantages when historical OD patterns are unavailable. We also conduct sensitivity analysis in terms of sensing error and snapshot frequency in imagery-derived density observations to evaluate the robustness of our model to various data availability conditions. The sensitivity analysis results indicate that the proposed framework is relatively insensitive to moderate systematic errors in density sensing, so that even with imperfect class-specific detection, the resulting OD estimates remain satisfactory under realistic levels of noise.
The primary use case of this study is calibration of city-wide dynamic network models, which serve as a testbed that reproduces status quo traffic states to support scenario analysis and policy evaluation. Especially when fixed detectors are sparse, unevenly deployed, or difficult to expand, satellite imagery with broader spatial coverage can be a promising data source used in developing these network models. From the data perspective, satellite images used in this study are infrequent, so our experiment is a proof-of-concept study under limited data availability. When more city-wide satellite imagery data are available, more tests can be conducted to evaluate model robustness, especially in large-scale real-world applications. Moreover, the proposed model can also accommodate video-based data, which provides traffic information of higher resolution, but this requires more experiments with only high-frequency traffic density information. From a specific application perspective, satellite images only capture the conditions of curbside spaces and outdoor parking lots. Fusing other parking information, such as indoor parking and parking meter transactions, will enable more comprehensive parking choice modeling. From a data preprocessing perspective, when class-specific information is unreliable or not available, the framework can also accommodate aggregated density information with satisfactory performance, as demonstrated in a preliminary study by Liu et al. [2024]. The computer vision pipeline can be further enhanced for accurate vehicle classification and road segmentation for map matching, particularly for freight vehicles, and its systematic detection biases can be explicitly corrected to further reduce potential bias in truck OD estimation.
@@ Acknowledgment
This research is sponsored by Fujitsu Research of America. ChatGPT 5.2 was used to assist in improving the clarity, conciseness, and overall quality of the writing in this manuscript. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of this manuscript.
@@
$P^{rs}_{i}$ path set for vehicle class $i$ between rsrs
$f^{rs}_{k,i,t}$ flow on path $k$ for vehicle class $i$ departing at time $t$ between OD pair rsrs
$c^{rs}_{k,i,t}$ travel cost of path $k$ for vehicle class $i$ departing at time $t$ between OD pair rsrs
$p^{rs}_{k,i,t}$ route choice proportion of choosing path $k$ for vehicle class $i$ departing at time $t$ between OD pair rsrs